Brendi,Nicki sell memberships in an automobile club over the telephone. Brenda sells one membership every 0.125 hour and nicki sells one membership every 0.1 hour. At what rate ( in memberships per hour) are the memberships being sold when both are working?
Answer by Earlsdon(6103)

If Brenda sells 1 membership every 0.125 hour, then she sells 1/0.125 memberships in 1 hour. 1/0.125 = 8 memberships per hour.
If Nicki sells 1 membership every 0.1 hour, then she sells 1/0.1 memberships in 1 hour. 1/0.1 = 10 memberships per hour.
Together, Brenda & Nicki sell 8 + 10 = 18 memberships per hour.